### v1

# The Effect of Similarity Measures on Accurate Stability Estimates for Local Surrogate Models in Text-based Explainable AI

Abstract
Recent work has investigated the vulnerability of local surrogate methods to adversarial perturbations on a machine learning (ML) model’s inputs, where the explanation is manipulated while the meaning and structure of the original input remains similar under the complex model. While weaknesses across many methods have been shown to exist, the reasons behind why still remain little explored. Central to the concept of adversarial attacks on explainable AI (XAI) is the similarity measure used to calculate how one explanation differs from another A poor choice of similarity measure can result in erroneous conclusions on the efficacy of an XAI method. Too sensitive a measure results in exaggerated vulnerability, while too coarse understates its weakness. We investigate a variety of similarity measures designed for text-based ranked lists including Kendall’s Tau, Spearman’s Footrule and Rank-biased Overlap to determine how substantial changes in the type of measure or threshold of success affect the conclusions generated from common adversarial attack processes. Certain measures are found to be overly sensitive, resulting in erroneous estimates of stability.
The Effect of Similarity Measures on Accurate Stability Estimates for Local Surrogate Models in Text-based Explainable AI
Christopher Burger Dept of Computer Science University of Mississippi cburger@olemiss.edu                        Charles Walter Dept of Computer Science University of Mississippi cwwalter@olemiss.edu                        Thai le Dept of Computer Science Indiana University tle@iu.edu
1 Introduction
Continuous advancements in machine learning and AI have enabled the proliferation of powerful, complex models throughout much of our daily lives. These advanced models are used in many disparate fields but disciplines with especially high consequences for failure, like medicine or finance, must be even more concerned with model error.
All models are subject to errors, but, when failure occurs, how is blame to be assigned? Complicating this assignment of blame is the type of failure that occurs. A model built for bird classification that misidentifies an image of a pine warbler as a yellow-throated vireo likely has no appreciable cost of failure except possible embarrassment. A collision detection system that interprets a person standing in a crosswalk as only aberrant noise is an immeasurably greater failure.
But when any model fails, it is important that its stakeholders to be informed of why such failure occurs. With ever increasing model complexity, being able to directly interpret the actions of the model is no longer feasible. As a response to this complexity the discipline of explanatory AI (XAI) seeks to develop tools to allow both developers and users of models to understand why a model works the way it does. A common method in XAI is to generate an approximation for the complex model, one that is inherently interpretable. The output of these approximations, or surrogate models, are then used to understand why a model has made some prediction. Of the surrogate models, local surrogates are largely used due to their increased fidelity since local surrogates concern themselves only with explaining an individual prediction (or at most a small subset), and do not attempt to explain the model as whole. But as surrogate models are models, they too can be subject to the same issues as the models they attempt to explain.
1.1 XAI Stability
One such property is stability or robustness, where an insignificant change in input should result in a correspondingly insignificant change in output Alvarez-Melis and Jaakkola (2018). This property is leveraged in adversarial attacks, where a model that is unstable can return obviously erroneous results given small changes to the input. Local surrogate models have been shown to lack stability across many types of data, including image, tabular, and text Slack et al. (2020); Ivankay et al. (2022); Burger et al. (2023). That is, the explanation of the complex model can be appreciably different given an insignificant change to the input. If an XAI model cannot produce consistent explanations across very similar documents when the output of the complex model itself remains similar, then the explanations from the XAI method become suspect. Suspect explanations cannot be trusted, and so the original complex model still remains opaque and may be prevented from use due to legal or social ramifications despite superior efficacy over other models.
Original Explanation Perturbed Explanation
Feature Weight Feature Weight
1 heartburn 1.77 1 heartburn 2.16
2 eat 0.59 2 choking 0.35
3 vomit 0.35 3 puked 0.34
4 choking 0.26 4 eat 0.33
5 feel 0.17 5 like 0.11
6 lot 0.15 6 pain 0.09
Original Text
I have a lot of heartburn and I feel like I’m choking when I eat. I also have a lot of stomach pain and I vomit a lot.
Perturbed Text
I have a lots of heartburn and I feel like I’m choking when I eat. I also have a lot of stomach pain and I puked a lot.
Comparative Similarities
RBO0.7 RBO0.9 Jaccard Jaccardw
0.69 0.74 0.72 0.90
Kendall Kendallw Spearman Spearmanw
0.10 0.48 0.50 0.50
Table 1: Perturbations generated under similarity measure RBO0.5 with final similarity of 0.75
1.2 Similarity Measures in XAI
Previous work in XAI stability has focused on the existence of instability given some fixed maximum amount of change that is applied to the input. This contrasts to standard adversarial attacks on (for example) classification models. These attacks often also have some perturbation limit, but the process has a clearly defined end point which is (if successful) a change in the predicted class. Prior work in XAI stability generally uses a similar approach to that of standard adversarial attacks, but lacks this clearly defined metric of success, and instead terminates at search exhaustion or perturbation limit Burger et al. (2023).
Original 26.5% Sim. 28.9% Sim.
Feature Weight Feature Weight Feature Weight
1 dogs 2.60 tennis 3.16 tennis 3.04
2 balls 2.10 dogs 2.65 dogs 2.68
3 helpful 0.44 balls 1.61 balls 2.22
4 fetching 0.39 adore 0.69 love 0.47
5 love 0.29 useful 0.49 fetching 0.42
6 play 0.07 toy 0.34 toy 0.39
7 tennis 0.04 wish 0.03 helpful 0.39
8 wish 0.03 fetches 0.02 wish 0.09
Original Text
i love dogs ! though i wish mine was more helpful while i play tennis . fetching balls . . .
Perturbed Text - 26.5% Similarity
i adore dogs ! though i wish mine was more useful while i toy tennis . fetches balls . . .
Perturbed Text - 28.9% Similarity
i love dogs ! though i wish mine was more helpful while i toy tennis . fetching balls . . .
Table 2: Perturbed explanations with close similarity values (calculated with respect to similarity measure RBO0.5) despite substantial differences in perturbation rate.
As the existence of instability has already been demonstrated, we seek to refine this knowledge by asking “what is the importance of the similarity measure used to guide the adversarial search process?”. The similarity measure is the engine for selection of appropriate perturbations in the search process and directly controls the acceptance of perturbations and the overall determination of success or failure of the algorithm. In particular we ask: Are certain similarity measures superior to others in providing quality adversarial explanations or accurate results in terms of XAI robustness? If so, under what conditions? Specifically, how do different measures compare in terms of sensitivity, or propensity to show a difference in similarity with respect to another measure. In Table 1 a perturbed explanation with 75% similarity to the original calculated using measure RBO0.5 (Section 3.2) is compared with eight other measures, resulting in up to 65% lower similarity between documents solely from choosing a new measure.
Original Text
“I have been having headaches for a while now. They are usually on the left side of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I have tried taking over-the-counter pain relievers, but they don’t seem to help much.”
Word Replacements Required to Satisfy Threshold
60% Threshold
Original Perturbed
pain $\implies$ agony
50% Threshold
Original Perturbed
pain $\implies$ agony
help $\implies$ assistance
40% Threshold
Original Perturbed
pain $\implies$ agony
seem $\implies$ appears
help $\implies$ aiding
tried $\implies$ strived
30% Threshold
Original Perturbed
pain $\implies$ agony
seem $\implies$ appears
help $\implies$ aiding
tried $\implies$ strived
having $\implies$ assuming
side $\implies$ sides
have $\implies$ ai
Table 3: Word Replacements required to satisfy each threshold. More stringent threshold generally necessitate more perturbations which reduces overall textual quality.
1.3 Minimal Perturbations
Additionally, does the choice of similarity measure affect the generation of minimum viable perturbations? Where given a threshold of success, what are fewest possible perturbations needed to reduce the similarity below this threshold while maintaining the structure and meaning of the original document. Clearly a method that consistently lacks stability when exposed to only a small number of perturbations is worse than one that that requires a consistently large subset of the text to be changed. In Table 2 we see the first document (26.5% similarity) requires substantially more perturbations to achieve a comparable level of similarity to that of the second attack (28.9% similarity). The presence of a single perturbation that provides a significant decrease in similarity implies a much more serious instability when compared to a method that requires multiple perturbations. Reducing the number of perturbations also helps maintain the quality of the perturbed text as ideal word replacements are difficult to generate given the vast search space. Table 3 provides an example of the process of quality degradation given repeated perturbation with the expanded text provided in Appendix A.
1.4 Restriction to Natural Language
We will restrict our focus to XAI models for text-based data. Our reasons are as follows: (1) Text-based models are ubiquitous as natural language is a primary communication mechanism. (2) As natural language is woven into the fabric of our daily experiences we as humans have an inherently robust sense for understanding the meaning of a document and so can judge the quality of a text document’s explanation effectively without any formal procedure or training. This innate ability to judge the quality of an explanation allows us to choose reasonable thresholds, reducing some of what is an inherently subjective decision. (3) We can appreciably simplify the comparison process between similarity measures. Restricting ourselves to only a single data type avoids ambiguity when comparing results from a given similarity measure between types of data. Additionally we can constrain the number of measures used, allowing use of those that may be more effective on text, but less transferable to other data. (4) Our depth of investigation can be increased significantly. Testing XAI stability is a computationally intensive process. Thus, focusing on one data type allows a much broader range of thresholds and similarity measures to be tested given a fixed amount of time and computational resources.
1.5 Contributions
Our contributions are summarized as follows.
  1. 1.
    Exploring the effect on the quality of stability estimates using different similarity measures to guide the adversarial search process in text-based XAI. Moreover, identifying measures that are unsuitable for use due to excessive sensitivity resulting in exaggerated indications of instability.
  2. 2.
    Determining if the choice of similarity measure has an impact on the number of perturbations to determine a successful attack. This allows greater discernment between XAI methods and their comparative levels of stability.
2 Background & Related Work
Prior work on XAI stability has emphasized on evaluating models using tabular or image data across various interpretation methods, which often use small perturbations to the input data to generate appreciably different explanations Alvarez-Melis and Jaakkola (2018); Ghorbani et al. (2017); Alvarez-Melis and Jaakkola (2018), or generate explanations that consist of arbitrary features Slack et al. (2020). Garreau et al. showed that key features can be omitted from the resulting explanations by changing parameters and that artifacts of the explanation generation process could produce misleading explanations Garreau and von Luxburg (2020, 2022). This was further extended to the analysis to text data Mardaoui and Garreau (2021) but only with respect to fidelity instead of stability of surrogate models.
Our work here is restricted to the least explored domain, text. Prior work exists directly involving adversarial perturbations for XAI but has been focused on determining the existence of such perturbations rather than establishing which components in the XAI method are most vulnerable Sinha et al. (2021); Ivankay et al. (2022); Burger et al. (2023). Other relevant work in text domain includes  Ivankay et al. (2022), which utilized a gradient-based approach but assumed white-box access to the target model; and Sinha et al. (2021), which generated adversarial attacks against black-box XAI methods. However their experiment design may have led to an underestimation of stability as explored in Burger et al. (2023) which investigated the inherent instability of the method of choice’s (Lime) sampling process for text data as well as provided an alternate search strategy focused on the preferential perturbation of features deemed unimportant.
2.1 XAI Method Selection
While there are many XAI methods available, we narrow our choices by selecting for three important criteria: proven usage in critical applications, a level of generalizability to other XAI methods, and the explanations generated are to satisfy certain attributes that constitute an effective explanation. Namely concision, order, and weight. Burger et al. (2023) From these criteria we choose Local Interpretable Model-agnostic Explanations (LimeRibeiro et al. (2016) as our target explanatory algorithm Lime is a commonly used and referenced tool in XAI frameworks, which has been integrated into critical ML applications such as finance Gramegna and Giudici (2021) and healthcare Kumarakulasinghe et al. (2020); Fuhrman et al. (2022). To explain a prediction, Lime trains a shallow, inherently explainable surrogate model such as Logistic Regression on training examples that are synthesized within a neighborhood of an individual prediction. The resulting explanation is an ordered collection of features and their weights from this surrogate model that satisfies our requirements for a quality explanation outlined above. For text data, explanations generated by Lime have features that are individual words contained within the document to be explained, which can be easily understood even by non-specialists. We note that our method is applicable to any XAI method that returns an explanation in this format, satisfying our criteria of generalizability.
We note that as our goal is to investigate more precisely the importance of the similarity measure used to guide the adversarial process Lime is not the focus of our inquiry but used as a familiar standard due to its substantial base of prior work and common use. As the prior work considers only if the perturbed input’s explanation is sufficiently different at the end of a process, often consisting of many perturbations, we instead ask, what would be reasonable similarity thresholds for stopping this process early?
3 Problem Formulation
As our goal is fundamentally similar to prior work in that we seek the discovery of perturbations that induce instability in XAI methods our constraints when generating adversarial perturbations are generally equivalent to the prior work that has established the existence of instability (in particular Lime. Namely, the meaning of the input is retained, as well as an identical predicted class for the perturbed input under the original model. We use the general perturbation process outlined in Burger et al. (2023) with one addition. While we retain a maximum amount of perturbations, we also use a threshold of similarity, $\tau$, used to signify a successful attack.
3.1 Search Procedure
The general search process is focused on the comparison of explanations, and not locating appropriate perturbations. We use the greedy search procedure standard to previous work where the indices of the words within the original document are ordered by importance (or lack thereof) to the model to be explained. The importance is calculated by removing the word at the $i^{th}$ index and determining the change in predicted probability. These indices are sorted, filtered by the constraints, and then iterated through where the word at index $i$ has perturbations generated to replace it. These perturbations are the the $n$ nearest neighbors in some embedding space where the final replacement is chosen by largest similarity decrease.
3.2 Similarity Measures
We assume the explanations generated are ranked lists, ordered by importance to the surrogate model (as is standard for Lime and common with other XAI methods). It is natural to then choose measures for similarity or distance designed specifically for ranked lists. From the candidates we select four popular measures that represent two major focuses of comparison, set based overlap and dissonance between paired features. For the following definitions, let $A$ and $B$ be ranked lists composed of unique features.
For measures that use sets as a primary component we choose the ubiquitous Jaccard index, and Rank-biased Overlap. The Jaccard index Eq. (1) is simply the ratio of the size of the intersection over the size of the union for two given sets. Intuitive and inherently bounded within $[0,1]$, the Jaccard index provides a useful baseline to compare against other measures. Its simplicity however renders it fairly coarse in determining similarity; For $Jaccard(A,B)=0.50$, a similarity of $50\%$, this requires at least $\frac{1}{3}$ of the words in B to have been perturbed (to a word not already in $A$). This is beyond the capability for current XAI perturbation methods to maintain the context of the original text consistently. However, the inherent instability (the difference in identical documents with altered initialization parameters) within Lime can at times be substantial, enough to render the measure sensitive to what is ultimately “natural” variation. This sensitivity can be reduced by applying weights to the elements, an extension we include for comparison.
$Jaccard(A,B)=\frac{|A\cap B|}{|A\cup B|}$ (1)
Rank-biased Overlap (RBO) Webber et al. (2010) is a summation of successively larger intersections, with each weighted by a term in a convergent series. This weighting scheme is controlled by a parameter $p\in(0,1)$ that can be adjusted to ascribe more or less weight to the top-k features, a useful property for explanations as concision is an important factor of explanability. Unbounding the surrogate model’s number of features results in many features of approximately zero weight, which for regression models often chosen as the surrogate this is effectively a judgment that the feature has no importance to the model being explained. While RBO was formulated to take in lists of arbitrarily large length Eq. 2 displays the version restricted to a list of known size. Here the size, $d=max(|A|,|B|)$.
$\begin{gathered}RBO(A,B,p,d)=\\ \frac{|A_{:d}\cap B_{:d}|}{d}p^{d}+\frac{(1-p)}{p}*\sum_{i=1}^{d}p^{i}\frac{|A_{:i}\cap B_{:i}|}{i}\end{gathered}$ (2)
For the measures of paired features we choose Kendall’s Tau Rank Distance and Spearman’s footrule. Both measures are in common use and are closely related, each a specific instance of a general correlation coefficient. Generally both measures return a distance rather than a similarity, but can be easily converted as both have a simple maximum bound described below.
Kendall’s Tau counts the number of pairwise inversions between $A$ and $B$ ( Eq.3) where $\mathbbm{1}[\cdot]$ is the indicator function.
$\sum_{i=1}^{max(|A|,|B|)}\mathbbm{1}[A[i]\neq B[i]]$ (3)
The above formulation is an extension with the capability of handling lists that are not of equal size. For such lists if $||A|-|B||\neq 0$ this value is added to the summation in Eq. (3). Clearly the maximum possible amount of dissonant pairs is $max(|A|,|B|)$ which provides the denominator when converting Eq. (3) to similarity. Other measures, especially those that calculate distance, may not have an upper bound. For our analysis we restrict ourselves to measures that have a bounded maximum distance, and so can be readily converted to a similarity value in $[0,1]$.
Spearman’s footrule (Eq. 4) is the sum of the difference between the location $i$ of each feature $a\in\mathbf{A}$ to its corresponding location $j$ in $\mathbf{B}$.
$\sum_{a\in\mathbf{A}}|i-j|$ (4)
Spearman’s footrule is effectively the $l_{1}$ distance applied to ranked lists. Similar to Kendall’s Tau, the footrule is by default not suited for disjoint lists, but the footrule distance is bounded, with maximum total distance of $\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor$ with an individual element having at most $|\mathbf{A}|-1$ possible distance. This leads to a natural choice for a penalty value for missing elements of $\frac{|\mathbf{A}|}{2}$, as for two completely disjoint lists we have $\sum_{i=1}^{|\mathbf{A}|}\frac{|\mathbf{A}|}{2}=\frac{|\mathbf{A}|^{2}}{2}$. Using a penalty the total maximum footrule distance can increase (Appendix B) but we note that this adjustment is not required as the proportion of disjoint elements between explanations is often small. As such, the original bound still remains useful even without the added penalty factor, which may avoid concerns of an arbitrary or ill-justified choice penalty. The weighted implementation for the footrule renders this concern even less meaningful as the disjoint features within the explanation are concentrated within features of low importance and so are assigned very little weight. We note that the extension to disjoint lists can induce asymmetry depending on the construction of the measure. Our work calculates the adjusted footrule value dependent on the original explanation being compared to perturbed.
Except for RBO which has inherent weighting, all of the previous measures have been seen in their unweighted format. The definitions are similar in structure and we refer the reader to Sculley (2007) Kumar and Vassilvitskii (2010) for background and their derivations. For our experiments we apply the same weights to each measure, which are the weights associated with the original unperturbed explanation which are normalized with respect to the absolute value of the weight. For example, for features $a,b$ with respective weights $0.25,0.10$ then changing $a$ and $b$ equally under our measure (in terms of distance, dissonance, etc) results in $a$ affecting the similarity more than $b$. These weighted measures are denoted using the subscript w.
3.3 Success Thresholds
We return to the question of where to stop the adversarial perturbation algorithm to balance fidelity with the original document and a sufficiently different explanation. As what constitutes a sufficiently different explanation is subjective, we instead select a range of values across each of the similarity measures to provide a broad comparison on common choices and to demonstrate the resulting quality of the perturbed document when a larger divergence from the original explanation is required. For our success threshold $\tau$, we choose levels of 30%, 40%, 50%, and 60% for algorithm termination. Thresholds near 70% begin to approach the inherent variation levels of LIME for certain datasets or non-standard sampling rates Burger et al. (2023). For thresholds lower than 30%, the greedy search process is generally too inefficient to find appropriate perturbations for most similarity measures within the maximum perturbation limitations imposed. For measures that require a weighting parameter (RBO) we use the following values 0.5, 0.7, and 0.9. For this formulation values approaching 0 increase the importance of the top features over the remaining features. Values close to 1.0 provide a less concentrated distribution of weight.
$\tau$ RBO0.5 RBO0.7 RBO0.9 Jaccard Jaccardw Kendall Kendallw Spearman Spearmanw
30% 0.12 0 0 0.02 0 0.95 0.52 0.14 0.02
GB
40% 0.26 0.07 0 0.24 0 0.98 0.64 0.38 0.12
50% 0.40 0.29 0.07 0.88 0 1 0.74 0.83 0.43
60% 0.40 0.40 0.28 1 0.05 1 0.81 1 0.69
30% 0.06 0.02 0 0.06 0 1 0.42 0.18 0.04
S2D
40% 0.18 0.04 0 0.52 0 1 0.48 0.58 0.28
50% 0.24 0.2 0.08 0.98 0.02 1 0.72 0.92 0.60
60% 0.24 0.3 0.28 1 0.14 1 0.84 1 0.94
Table 4: Attack Success Rates on Lime under DistilBERT.
$\tau$ RBO0.5 RBO0.7 RBO0.9 Jaccard Jaccardw Kendall Kendallw Spearman Spearmanw
$\mu$ M $\mu$ M $\mu$ M $\mu$ M $\mu$ M $\mu$ M $\mu$ M $\mu$ M $\mu$ M
30% 0.28 0.29 - - - - 0.27 0.27 - - 0.19 0.22 0.14 0.15 0.26 0.28 0.28 0.28
GB
40% 0.34 0.32 0.38 0.39 - - 0.39 0.39 - - 0.26 0.29 0.19 0.24 0.35 0.36 0.37 0.38
50% 0.40 0.43 0.48 0.48 0.46 0.47 0.47 0.46 - - 0.32 0.32 0.28 0.33 0.47 0.48 0.47 0.47
60% 0.40 0.43 0.52 0.53 0.57 0.58 0.55 0.55 0.59 0.59 0.37 0.40 0.33 0.36 0.56 0.56 0.54 0.55
30% 0.27 0.28 0.28 0.28 - - 0.29 0.29 - - 0.22 0.23 0.19 0.21 0.26 0.28 0.25 0.25
S2D
40% 0.35 0.35 0.40 0.39 - - 0.37 0.38 - - 0.29 0.29 0.28 0.33 0.37 0.38 0.37 0.37
50% 0.42 0.46 0.48 0.49 0.49 0.49 0.47 0.48 0.47 0.47 0.36 0.38 0.39 0.44 0.47 0.48 0.45 0.46
60% 0.42 0.46 0.55 0.58 0.58 0.58 0.56 0.57 0.56 0.57 0.38 0.39 0.49 0.51 0.56 0.57 0.57 0.58
Table 5: Mean and Median Similarities of Successful Attacks.
4 Experiment Setup
To provide the raw material for our analysis, we generate batches of 50 adversarial examples using the algorithm in Burger et al. (2023). Each batch is generated with respect to a similarity measure and a success threshold. Nine similarity measures are used: The Jaccard index, Kendall’s Tau Rank Distance, and Spearman’s footrule, each in their standard and weighted implementations and RBO, with weighting parameters 0.5, 0.7, and 0.9. Combined with the thresholds of 30%, 40%, 50%, and 60% this results in 1,800 adversarial examples per dataset. For the datasets we use two of those included in Burger et al. (2023) and their associated pretrained models. The first being the short length (average of 11 words) gender bias Twitter dataset (GB) Dinan et al. (2020) and the second being the moderate length (average of 29 words) symptoms to diagnosis dataset (S2D) (Kaggle). The final and longest length dataset, IMDB movie reviews, proved computationally infeasible due to the excessive time requirements to generate the examples, approximately 125 days of continuous computation on a single A6000. The model to be explained is a DistilBERT Sanh et al. (2019) fine-tuned on each respective dataset. Our interests concern the generalized properties of the similarity measures and not their specific performance associated with a given XAI method and base model and so we choose for computational convenience a single efficient model to explain.
$\tau$ RBO0.5 RBO0.7 RBO0.9 Jaccard Jaccardw Kendall Kendallw Spearman Spearmanw
30% 0.14 - - 0.25 - 0.12 0.10 0.21 0.13
GB
40% 0.14 0.19 - 0.23 - 0.10 0.10 0.21 0.20
50% 0.07 0.15 0.25 0.22 - 0.09 0.09 0.18 0.17
60% 0.07 0.10 0.21 0.17 0.22 0.08 0.09 0.15 0.15
30% 0.11 0.2 - 0.23 - 0.06 0.08 0.17 0.16
S2D
40% 0.12 0.12 - 0.20 - 0.05 0.07 0.17 0.15
50% 0.05 0.15 0.19 0.17 0.21 0.04 0.07 0.14 0.15
60% 0.05 0.09 0.17 0.13 0.20 0.04 0.05 0.10 0.13
Table 6: Perturbed Word Rates for Successful Attacks
$\tau$ RBO0.5 RBO0.7 RBO0.9 Jaccard Jaccardw Kendall Kendallw Spearman Spearmanw
USE PPL USE PPL USE PPL USE PPL USE PPL USE PPL USE PPL USE PPL USE PPL
30% 0.86 1.10 0 0 0 0 0.84 5.06 0 0 0.87 0.94 0.87 0.75 0.82 3.61 0.84 1.59
GB
40% 0.85 2.04 0.84 2.34 0 0 0.82 3.76 0 0 0.88 0.68 0.88 0.66 0.82 1.79 0.81 1.77
50% 0.89 0.51 0.84 1.98 0.80 5.66 0.83 2.25 0 0 0.89 0.65 0.88 0.61 0.84 1.88 0.84 1.52
60% 0.89 0.51 0.87 0.89 0.81 2.12 0.85 1.60 0.82 4.51 0.90 0.45 0.89 0.50 0.85 1.26 0.85 1.55
30% 0.90 1.46 0.82 3.59 0 0 0.83 6.35 0 0 0.93 0.78 0.90 1.26 0.86 3.04 0.90 2.62
S2D
40% 0.89 1.69 0.89 2.02 0 0 0.84 4.02 0 0 0.94 0.48 0.92 0.89 0.86 3.06 0.87 2.01
50% 0.92 0.62 0.88 2.21 0.85 3.63 0.86 3.03 0.80 5.10 0.94 0.36 0.92 0.91 0.87 1.91 0.87 2.35
60% 0.92 0.62 0.90 1.20 0.86 3.68 0.88 1.98 0.82 3.23 0.95 0.30 0.93 0.67 0.90 1.32 0.88 1.75
Table 7: Perturbed Document Quality - USE is the Universal Sentence Encoder, PPL is perplexity.
5 Results and Discussion
All values in the following tables are calculated only for successful attacks. The symbol  -  indicates non-applicability due to no successful attacks on that particular combination of similarity measure, threshold, and dataset.
Attack Success Rate (Table 4): Immediately seen is the inappropriateness of certain similarity measures for use in text-based adversarial XAI attacks. Kendall’s tau is extremely sensitive with nearly 100% attack success across every combination of threshold and dataset. This level of sensitivity renders the measure useless here as the inherent instability of Lime is the cause, little perturbation induced instability is present. This follows for the weighted version as well, though the success rates are not quite as high. Jaccard, Spearman, and Spearmanw also show excessive sensitivity for the higher similarity thresholds of 50% and 60%.
Jaccardw and RBO0.9 exhibit the opposite behavior instead being coarse with few, if any, successful attacks under all but the most lenient threshold. This is not inherently negative, as both measures effectively require the top few features to change, with complete removal from the explanation in the case of Jaccardw and at least a substantial decrease in ranking for RBO0.9. Since much of the weight of an explanation is often associated with the top features, these measures may prove useful as a fast heuristic for a substantial explanation difference.
The attack success rates are generally consistent between the datasets. As the sampling rate for Lime was optimized according to prior work, the inherent instability of the sampling process was kept similar despite the size difference.
Perturbation Quality (Tables 6,7): The quality of the perturbed document follows the expected pattern of more perturbations results in lower quality. We exclude the Kendall measures from our discussion here as they are extremely sensitive and so few perturbations are required to meet the required threshold. Our measures for quality are perplexity (PPL) and similarity as calculated using cosine similarity with the universal sentence encoder (USE).
The coarseness of Jaccardw and RBO0.9 results in the worst document quality with both the lowest similarity and the highest perplexity. Jaccard also has poor document quality. Both Spearman and Spearmanw provide comparable results for the levels of perturbation given document quality. The RBO variants follow the expected trade-off of increased perturbation rates with subsequently reduced quality as the measure becomes coarser.
Minimal Perturbations: There are no strong indications for the choice in similarity measure being a major factor in finding minimal perturbations. We see no average ending similarities appreciably below the success threshold (with the exception of Kendall due to its sensitivity) (Table 5) and there are no notable signs of consistently low perturbation rates (Table 6). However, the search process is very greedy, taking the first viable perturbation that reduces the explanation similarity. An alternative search strategy would likely be superior when focused exclusively on minimal perturbations. Either with a threshold used to specify a minimum reduction in similarity in accepting a perturbation or an entirely new algorithm that can explore the search space more effectively.
Overall: The Kendall and Jaccard measures are of limited use due to sensitivity or poor perturbed document quality. RBO in general provides a good balance between sensitivity and perturbed document quality across all thresholds at the cost of requiring manual adjustment of the weighting parameter for the given threshold. The Spearman measures show promise in comparison with RBO for more demanding thresholds but suffer from sensitivity for larger values. Spearmanw may prove particularly useful if the original explanatory weights are deemed important or the fine-tuning of RBO too cumbersome. No particular measure stands out as substantially affecting the search procedure in terms of search efficiency or locating a set of minimal perturbations for a given threshold.
6 Conclusion
We have shown that a poor choice in similarity measure can drastically skew the results of an adversarial attack, either by selecting a measure that is too sensitive and so overstating the weakness of the XAI method, or by choosing a measure too coarse and overstating the method’s resilience. Practitioners and researchers should choose these measures judiciously, keeping in mind how sensitive a measure is to the inherent variation between explanations. Our findings confirm prior work in that measures that use both order and weight produce results with fewer perturbations and subsequently higher textual quality. However, no measure tested proved ideal in selecting for minimal perturbations. It appears necessary that the search process itself will also need adjustment as the current greedy approach is consistent regardless of measure used.
Limitations
Transferability to other XAI methods: While the fundamentals of the discussion on similarity measures is not unique to Lime, our general conclusions on the fitness of particular similarity measures may not be applicable on different XAI methods. In particular, methods with less inherent instability may see more useful results from measures like Kendall’s Tau where its excessive sensitivity and subsequently exaggerated attack effectiveness may no longer pose a problem.
Computational Expenditure: The computational resources needed to produce the adversarial examples was significant. Existing methods used to generate the attacks were not designed for multi-thousand collections of examples to be generated. Purpose built tools focused on efficiency would allow larger numbers of samples to be reasonably generated which may result in alternate conclusions.
Other similarity measures: The similarity measures used here were designed for the use of comparison of ranked lists. Other measures with this purpose or more generalized measures exist and our conclusions may not be reflective of these untested measures.
Broader Impacts and Ethics Statement
The authors anticipate there to be no reasonable cause to believe use of this work would result in harm, both direct or implicit. The authors disclaim any conflicts of interest pertaining to this work.
Appendix A Additional Experimental Results
Table 8: Perturbed Explanations at different attack success thresholds. Degradation in textual quality is apparent with continued perturbation. Italics within an explanation indicate a word subject to perturbation or its perturbed form.
Top-6 Features
Original
Feature Weight
1 headaches 2.80
2 sensitivity 0.30
3 having 0.26
4 tried 0.25
5 sound 0.25
6 vomiting 0.21
60% Threshold
Feature Weight
1 headaches 2.85
2 counter 0.30
3 sound 0.25
4 don’t 0.23
5 sensitivity 0.19
6 having 0.16
50% Threshold
Feature Weight
1 headaches 3.01
2 having 0.38
3 sensitivity 0.34
4 taking 0.24
5 vomiting 0.20
6 don’t 0.18
40% Threshold
Feature Weight
1 headaches 2.86
2 sensitivity 0.30
3 having 0.25
4 strived 0.20
5 counter 0.17
6 light 0.17
30% Threshold
Feature Weight
1 headaches 3.06
2 sensitivity 0.27
3 strived 0.26
4 left 0.25
5 painful 0.22
6 counter 0.21
New Locations for original Top-6 Features
Feature Weight
1 headaches 2.80
2 sensitivity 0.30
3 having 0.26
4 tried 0.25
5 sound 0.25
6 vomiting 0.21
Feature Weight
1 headaches 3.10
5 sensitivity 0.19
6 having 0.16
9 tried 0.14
3 sound 0.25
8 vomiting 0.15
Feature Weight
1 headaches 3.01
3 sensitivity 0.34
2 having 0.38
12 tried 0.26
13 sound 0.0.08
5 vomiting 0.20
Feature Weight
1 headaches 3.06
2 sensitivity 0.27
13 having 0.25
3 strived 0.20
16 sound 0.15
13 vomiting 0.10
Feature Weight
1 headaches 3.06
2 sensitivity 0.27
13 assuming 0.22
3 strived 0.26
10 sound 0.15
18 vomiting 0.04
Original Text
I have been having headaches for a while now. They are usually on the left side of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I have tried taking over-the-counter pain relievers, but they don’t seem to help much.
Perturbed - 60% Similarity
I have been having headaches for a while now. They are usually on the left side of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I have tried taking over-the-counter agony relievers, but they don’t seem to help much.
Perturbed - 50% Similarity
I have been having headaches for a while now. They are usually on the left side of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I have tried taking over-the-counter agony relievers, but they don’t seem to assistance much.
Perturbed - 40% Similarity
I have been having headaches for a while now. They are usually on the left side of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I have strived taking over-the-counter agony relievers, but they don’t appears to aiding much.
Perturbed - 30% Similarity
I have been assuming headaches for a while now. They are usually on the left sides of my head and are very painful. I also get nausea, vomiting, and sensitivity to light and sound. I ai strived taking over-the-counter agony relievers, but they don’t appears to aiding much.
Appendix B Bounding Penalized Spearman’s Footrule Maximum Distance
Recall that Spearman’s footrule possesses a maximum total distance (when the list is completely inverted) of $\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor$ for $\mathbf{|A|}=\mathbf{|B|},\;\mathbf{A}\cap\mathbf{B}=\mathbf{A}$ with an individual element having at most $|\mathbf{A}|-1$ possible distance.
Given a penalty, $p\geq 1$, for disjoint elements (i.e there exists $e\in\mathbf{A}\;s.t\;e\notin\mathbf{B}$), we may have a total possible distance that exceeds the non-adjusted footrule.
Let $\mathbf{A},\mathbf{B}$ be lists that are possibly disjoint and without loss of generality assume $\mathbf{|A|}\geq\mathbf{|B|}$.
As the largest possible individual distance is $|\mathbf{A}|-1$ assume $p\geq\mathbf{|A|}-1$. Then trivially the maximum distance is $p\mathbf{|A|}$.
Now assume $p<\mathbf{|A|}-1$. The individual distances for a completely inverted list follow the patterns of: $\mathbf{|A|}-1,\mathbf{|A|}-3,\mathbf{|A|}-5...1,1,...\mathbf{|A|}-5,\mathbf{|A|}-3,\mathbf{|A|}-1$ for $\mathbf{|A|}$ even, and $\mathbf{|A|}-1,\mathbf{|A|}-3,\mathbf{|A|}-5...0...\mathbf{|A|}-5,\mathbf{|A|}-3,\mathbf{|A|}-1$ for $\mathbf{|A|}$ odd.
We can view the new maximum penalty as the sum of two components: $p_{g}$, the individual distances greater or equal to the penalty, and $p_{l}$, those less than the penalty.
As $p<\mathbf{|A|}-1$ then there exists some distance $d\;s.t\;p\leq d\leq\mathbf{|A|}-1$. Let this $d$ be located at index $i$ within the pattern $\mathbf{|A|}-1,\mathbf{|A|}-3,\;...$ , where $\mathbf{|A|}-1$ is index $1$. The size of this subset of the above pattern is $\mathbf{|A|}-2i$ and so its sum can be given as $\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
Then $p_{g}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor-\lfloor\frac{({\mathbf{{|A|}}-2i)}^{2}}{2}\rfloor$. And since $p\geq d$ then $p_{l}=p(\mathbf{{|A|}}-2i)-\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.
So the maximum distance for $p<\mathbf{|A|}-1$ is given by $p_{g}+p_{l}=\lfloor\frac{|\mathbf{A}|^{2}}{2}\rfloor+p(\mathbf{{|A|}}-2i)-2\lfloor\frac{({|\mathbf{A}|-2i})^{2}}{2}\rfloor$.